RHIC Running: Nine hour store yesterday with polarization 55%/48% in B/Y. Vernier scan completed, manual cogging tested and it was found that Blue bucket 1 (bunch 1) was colliding with Yellow bucket 4 (bunch 2) at PHENIX during fill 9798. Two ramps failed due to no separation at JET. The second pulled the permit. Reverted to an old fill. 1.3x1.3e11 with 109x109. PHENIX ZDC was ~10kHz at start of fill. PASS dropped the next store after 4 hours. PASS tripped another gate, like last week.
APEX is running all day. Preliminary results are that the 60cm b*-squeeze is working well, so previous plans to take 4 hours tomorrow to try new ramps at intermediate b*. will not happen. We will go straight to 60cm. This will be attempted tomorrow during the day. At the end of APEX today we will switch the JET from Blue to Yellow.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |